OpenStudy (lgbasallote):

at 1:00 pm a thermometer reading 10F is removed from a freezer and placed in a room whose temperature is 65 F. at 1:05 pm, the thermometer reads 25 F. Later, the thermometer is placed back in the freezer. at 1:30 pm the thermometer reads 32 F. when was the thermometer returned to the freezer and what was the thermometer reading at that time? Answer: 1:20 pm; 50.15 F

OpenStudy (lgbasallote):

the formula is \[\huge \ln (\frac{T-T_m}{T_o - T_m}) = kt\]

OpenStudy (lgbasallote):

where T = final temp T_o = initial temp T_m = outside temp k = constant t= time

OpenStudy (lgbasallote):

can you do it @nbouscal ?

OpenStudy (anonymous):

How can the answer be that it was reading 50.15 F at 1:30 pm if the question says that it read 32 F at 1:30 pm?

OpenStudy (lgbasallote):

oops typo..that's 1:20

OpenStudy (lgbasallote):

the answer should be 1:20

OpenStudy (anonymous):

Hrm. No, I do not know how to solve this problem.

OpenStudy (lgbasallote):

aw darn :(

OpenStudy (lgbasallote):

@Limitless can you?

OpenStudy (anonymous):

\[k=\frac{\ln\left(\frac{T-T_m}{T_0-T_m}\right)}{t}\] \(T_0=10\), \(T=25\), \(T_m=65\), \(t=1:05-1:00=5\) \[k=\frac{\ln(40/55)}{5}=\frac{\ln(8/11)}{5}\] I'm stuck at reasoning out how to pinpoint the position in the gap of 1:05 to 1:30.

OpenStudy (lgbasallote):

i think i got an equation differen from this one?

OpenStudy (lgbasallote):

lol nevermind i have no idea what to do here

OpenStudy (lgbasallote):

so k = -0.0637?

OpenStudy (anonymous):

If my logic is correct and you used the appropriate calculator, yes.

OpenStudy (zarkon):

I don't get exactly what you have written as the answer

OpenStudy (zarkon):

i get 1:20.555 and a temp of 50.147

OpenStudy (anonymous):

@Zarkon, that's part of the answer... I think.

OpenStudy (anonymous):

I did not write a complete answer; I just posted my thoughts so far.

OpenStudy (zarkon):

Talking about answer in the OP

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

It's rounding, by the way.

OpenStudy (zarkon):

I made a couple assumptions to do it though

OpenStudy (zarkon):

I assumed the freezer temp was 10F. and so k was the same for growth and decay

OpenStudy (anonymous):

That is a valid assumption, I think.

OpenStudy (zarkon):

And I used my calculator to solve the equations :) too lazy to do by havnd

OpenStudy (zarkon):

*hand

OpenStudy (lgbasallote):

yes..i used calculator too..but i think i got 8 or something

OpenStudy (lgbasallote):

could you post the ln you got? i think that's where i went wrong

OpenStudy (zarkon):

so I get \[T(t)=65-55\left(\frac{8}{11}\right)^{t/5}\] my calculator simplified it to this....for the first part

OpenStudy (zarkon):

then I made the equation \[T_2(t,s,w)=10+(s-10)e^{-(t-w)\frac{\ln(11/8)}{5}}\] then I solved \[T_2(30,T(t),t)=32\] for \(t\) ;)

OpenStudy (zarkon):

got \[t=20.5554814377\]

OpenStudy (lgbasallote):

uhmm maybe i should post my solution and you can spot where i went wrong...

OpenStudy (lgbasallote):

because i cannot see my mistake from your solution since beginning is different

OpenStudy (zarkon):

and \[T(\text{the above})=50.1479302375\]

OpenStudy (zarkon):

my \(T(t)\) can be written as \[T(t)=65-55e^{-t\frac{\ln(11/8)}{5}}\]

OpenStudy (zarkon):

which I believe should be the same as Limitless

OpenStudy (zarkon):

I'm out of pepsi now so my powers might be fading

OpenStudy (lgbasallote):

ugh i dont understand this T_T

OpenStudy (zarkon):

I don't know if I can eloquently describe what I did

OpenStudy (zarkon):

we have two main issues here...we dont know the time it went back into the freezer and thus we dont know the temperature

OpenStudy (lgbasallote):

okay i set up two conditions first condition T_m is 65 second contition T_m = 10 is that right

OpenStudy (zarkon):

my \[T_2(t,s,w)\] has 3 variables...the first t is the time at the end which will be 30 the S is the temperature at time (t) (unknown...it is \(T(t)\) and w is the time at which we put it back in the frezer

OpenStudy (lgbasallote):

hmm i see

OpenStudy (lgbasallote):

t is the time before 30?

OpenStudy (zarkon):

we then know that \[T_2(30,T(t),t)=32\] since this is the temp at time 30

OpenStudy (zarkon):

yes

OpenStudy (lgbasallote):

wait...ill try to re-solve

OpenStudy (lgbasallote):

i quit..i cant do this...sorry...and thanks for all your help

OpenStudy (zarkon):

sorry I couldn't help you more